Suppose the European Union (EU) is investigating a proposed merger between two of the largest distillers of premium Scotch liquor. Based on some economists’ definition of the relevant market, the two firms proposing to merge enjoyed a combined market share of about two-thirds, while another firm essentially controlled the remaining share of the market. Additionally, suppose that the (wholesale) market elasticity of demand for Scotch liquor is -1.4 and that it costs $15.90 to produce and distribute each liter of Scotch.

Based only on these data, provide quantitative estimates of the likely pre- and postmerger prices in the wholesale market for premium Scotch liquor.

Instruction:Do not round intermediate calculations. Enter your final responses rounded to the nearest penny (two decimal places).

Pre-merger price: $____

Post-merger price: $_____
Answer #1

Use P = MC * (ed)/(ed + 1)

where ed is price elasticity and MC is marginal cost

Premerger condition has 3 firms. Hence pre-merger price is ( 3 x-1.4 P=15.90 (3x-1.4 +1) = 15.904 (132) = 20.87 Post-merger c